# ColorFoil: Investigating Color Blindness in Large Vision and Language Models

### v1

@@ 4 Experimental Setup
Table 1: Experiment results. We evaluate seven of the state-of-the-art V&L models on the MS COCO and Flickr30k subsets from ColorFoil. Accuracy (%) and F1-scores are reported. We conduct three experiments in which the models are presented different number of foils (modified caption) along with the original caption. The V&L models tend to struggle in challenging conditions with more foils. BridgeTower and ViLT outperform other V&L models including CLIP and its variants and GroupViT by a large margin.
@@
ALIGN [11] 86.03 0.93 87.70 0.93 79.47 0.89 81.43 0.90 71.03 0.83 74.57 0.85
AltCLIP [5] 83.41 0.92 82.69 0.91 77.18 0.87 73.68 0.85 69.08 0.82 64.39 0.78
BridgeTower [29] 97.26 0.99 96.83 0.98 95.71 0.98 94.46 0.97 92.61 0.96 90.81 0.95
CLIP [20] 84.37 0.92 85.24 0.92 76.82 0.87 76.26 0.87 67.37 0.81 68.33 0.81
CLIPSeg [16] 83.05 0.91 82.01 0.90 74.00 0.85 72.97 0.84 64.56 0.78 63.07 0.77
GroupViT [28] 82.73 0.91 81.64 0.90 73.10 0.84 71.77 0.84 63.80 0.78 62.12 0.77
ViLT [12] 95.53 0.98 94.29 0.97 92.83 0.96 91.85 0.96 88.74 0.94 87.38 0.93
@@ 7 Ethical Considerations
Ensuring reproducibility is a crucial aspect of scientific research. To foster open research practices, we have made our code publicly accessible, allowing others to reproduce and verify our findings. [^1] Code is available at https://github.com/samin9796/ColorFoil

### v2

@@ 4 Experimental Setup
Table 1: Experiment results. We evaluate seven of the state-of-the-art V&L models on the MS COCO and Flickr30k subsets from ColorFoil. Accuracy (%) and F1-scores (%) are reported. We conduct three experiments in which the models are presented different number of foils (modified caption) along with the original caption. The V&L models tend to struggle in challenging conditions with more foils. BridgeTower and ViLT outperform other V&L models including CLIP and its variants and GroupViT by a large margin.
@@
ALIGN [11] 86.03 93.32 87.70 93.45 79.47 88.75 81.43 89.76 71.03 83.06 74.57 85.43
AltCLIP [5] 84.89 91.82 82.69 90.52 77.29 87.19 73.68 84.84 69.08 81.71 64.39 78.34
BridgeTower [29] 97.31 98.63 96.83 98.32 95.71 97.81 94.46 97.15 92.61 96.16 90.81 95.18
CLIP [20] 84.42 91.55 85.24 92.07 76.19 86.49 76.26 86.53 67.37 80.50 68.33 81.18
CLIPSeg [16] 83.05 91.09 82.01 90.12 74.00 85.42 72.97 84.37 64.56 78.45 63.07 77.35
GroupViT [28] 82.73 91.67 81.64 89.89 73.10 83.98 71.77 83.57 63.80 77.89 62.12 76.63
ViLT [12] 95.69 97.79 94.29 97.06 92.83 96.28 91.85 95.35 88.74 94.04 87.38 93.27
@@ 7 Ethical Considerations
Ensuring reproducibility is a crucial aspect of scientific research. To foster open research practices, we will make our code publicly accessible, allowing others to reproduce and verify our findings.